# Tunneling method for Hawking quanta in analogue gravity

### v1

@@ I Introduction
Here, to address these shortcomings, we employ a quasi-local technique: the tunneling picture for Hawking radiation [20, 21]. This method uses a geometric optics approximation to describe the particle creation process as a quantum tunneling event, involving complex paths across the causal boundary. In contrast to the Bogoliubov coefficients, no knowledge of the system’s asymptotic boundary is necessary, allowing particle creation to be studied almost entirely locally, wherever the event may happen in spacetime. This property is especially useful for localising effects involving effective horizons, such as those associated with modified dispersion relations that characterise fields in analogue spacetimes.
@@ II General setting
[…]
In (1), the sign of $v(x)$ determines the nature of this spacetime. This line-element describes the Painlevé-Gullstrand coordinates of a sonic black hole if $v(x)\leq 0$ and of a sonic white hole if $v(x)\geq 0$ [23]. In the following we are going to focus on the former case. Hereinafter, we will omit to write the $x-$dependence explicitly, unless necessary. This geometry enjoys a Killing vector, associated to time translation invariance in the lab frame, given by $\chi^{a}\partial_{a}=\partial_{t}$.
Together with the lab frame, it will be useful to introduce the so-called preferred frame, which is the system of reference where the fluid is at rest, described by the two normalised vectors
@@ II.1 Perturbations with modified dispersion relation
[…]
This setup accommodates two free functions: the fluid’s velocity profile $v(x)$ and the dispersion relation function $F(k^{2},\Lambda)$. For what concerns our stationary flow’s velocity profile $v(x)$, we shall assume that this quantity can in general be described by a monotonous function and we will treat setups in which the flow admits a subsonic and a supersonic region (super-critical flow) or a purely subsonic setup (sub-critical flow) seperately. [^1] Later on, we shall also discuss purely supersonic flows, but only as a dual case w.r.t. to the purely subsonic ones. We address the limiting case in which the flow is everywhere sub-critical except for a sonic point (critical flow) too.
[…]
@@
[…]
@@ II.2 Particles
The constant phase contours of $\phi$ yield the trajectory of the associated ray. Introducing the four-momentum $k_{a}$ enables us to rewrite the field equation as a dispersion relation for a point particle, at the leading order in the WKB formalism
[…]
@@
Since $\Omega$ is associated to a translational symmetry, the idea is to find mode solutions $\phi_{\Omega}$ of (7) at fixed Killing energy, which can be proven to be a conserved quantity even for a modified dispersion [26]. This, together with the dependence of our modified dispersion relation on the effective field theory scale $\Lambda$, strongly suggest the introduction of a dimensionless parameter $\alpha\equiv\Omega/\Lambda$ that controls the deviations from the relativistic behaviour (recovered in the limit $\alpha\to 0$).
@@ II.3 Particle trajectories on the background
The boundary between these two regions depends on the energy and is located at the point $x_{\textsc{tp}}(\alpha)$ (with TP standing for “turning point”) where two out of the four solutions become degenerate. In terms of trajectories, this represents two smoothly merging trajectories at $x_{\textsc{tp}}(\alpha)$. The name turning point can be understood graphically in the $(t,x)$-plane. The curves at the meeting point can always be interpreted as two branches of a single trajectory that, in the $(t,x)$-plane, turns back at $x_{\textsc{tp}}(\alpha)$. Additionally, the position of $x_{\textsc{tp}}(\alpha)$ depends crucially on the sign of $\xi$. In the superluminal case, $x_{\textsc{tp}}(\alpha)$ can always be found inside the horizon, where $|\chi|^{2}>0$, while in the subluminal case $x_{\textsc{tp}}(\alpha)$ lies outside. This follows directly from the discriminant of (12). The shape of the trajectories is sketched for both cases in Fig. 1.
Figure 1: Example of the four different solution of (12) at fixed $\Omega$. The horizontal axis shows $v$, without specifying any profile $v(x)$ yet, while $c_{s}=1$ everywhere. The dashed black vertical line is the Killing horizon $v=-1$ in both figures. In all plots, we have taken $\alpha=0.02$ for each ray. Left panel: subluminal case ($\xi=-1$); we see that the turning point, where the dashed and the solid orange lines meet, lies outside the Killing horizon. Right panel: superluminal case ($\xi=1$) for which we find turning point to be inside the Killing horizon. Both cases share a regular mode (in blue) which travels inwards and a mode (in red) which lingers at the horizon. The latter changes its direction depending on the sign of $\xi$, while the blue one remains qualitatively unchanged.
@@ III Hawking radiation by tunnel method
In relativistic settings, i.e.  when considering phononic dispersion relations, the presence of a sonic/Killing horizon is sufficient for sonic black holes to emit particles. Essentially, such a process reflects the analogue of the standard Hawking effect [28]. Within a tunneling approach [20, 21], this correspondence can be easily seen by analysing the outgoing particle peeling off outwards from the horizon and its counterpart peeling off inwards on the other side. If $x_{\textsc{kh}}$ be the horizon’s location, the tunneling rate will be given by [29, 30]
@@
Mathematically speaking, in the above relativistic treatment, non-zero transition rates presume the presence of a simple pole in the expression of $\frac{\mathrm{d}t}{\mathrm{d}x}$. However, in a non relativistic scenario, nothing comparable will ever occur, given that the modified dispersion relation, never produces an exact peeling (see Fig. 1) and therefore the particle trajectories at the Killing horizon remain analytic. In turn, one would expect an absence of Hawking radiation in this settings (or at the least that this effect cannot be described via the tunneling method).
@@
In the following section, we shall focus on how to describe this effect in a simple, analytical way via the tunneling method. As we shall see, the extension of the latter to setups with modified dispersion relations will prove itself useful and informative, thus, nicely complementary to previous, semi-analytical or numerical studies  [12, 13, 14, 15, 16, 17, 18, 35, 19, 33], based on the Bogoliubov coefficients analysis.
III.1 Non-relativistic scenario: $\alpha\to 0$ limit
As already stressed, when $\alpha\to 0$, the upper branch of the turning mode on one side of the Killing horizon and the lingering mode on the other side tend to become the relativistic Hawking quanta. More specifically, these two solutions can be treated perturbatively [27] in $\alpha$ and their group velocity can be written as $c_{g}=1+\delta c_{g}(x,\alpha)+\dots$, with $\delta c_{g}(x,\alpha)\ll 1$. In the case of a dispersion relation like (7), $\delta c_{g}(x,\alpha)=\alpha^{2}f(x)$, with $f(x)$ depending on the geometry and on the sign of $\xi$. This allows us to write
$\frac{\mathrm{d}t}{\mathrm{d}x}=\frac{1}{c_{g}+v}=\frac{1}{1+v}-\frac{\alpha^{2}f(x)}{(1+v)^{2}}+\cdots,$ (16)
that causes a non-zero contribution when plugged into ${\rm Im}(S)$ as in Eq. (13), thus giving a non-trivial tunneling amplitude. The associated “effective temperature” (as it stems per se from an approximated trajectory) will then receive corrections of order ${\alpha}$ so that at the lowest order one gets
$T(\alpha)=\frac{\kappa_{\textsc{kh}}}{2\pi}\left[1+\alpha^{2}{\rm Res}\left(\frac{f(x)}{(1+v(x))^{2}},\,x_{\rm KH}\right)+\cdots\right]\,.$ (17)
This calculation unveils the capability of the tunneling formalism to describe the particle production phenomenon. Additionally, it becomes clear that the main protagonists in this effect are given by the upper branch of the turning mode and the lingering mode. Let us see now, how we can generalise our findings beyond the perturbative treatment around the $\alpha=0$ limit.
@@ IV Particle production: supercritical flow
Our starting point will be the supercritical flow, that is to say, the permitted range for $v(x)$ supports the presence of a horizon, i.e. $|v(x)|>1$ for some $x$. For now, we keep our treatment as general as possible without particularising to a specific profile for $v(x)$. We demand: i) $v(x)$ to be monotonous to avoid inner Killing horizons (namely, $v(x)=-1$ has a single root), ii) $v\to 0$ for $x\to-\infty$, and iii) conditions (20) and (23) hold almost everywhere apart from a small region around the effective horizon. A typical example of such flow is the one considered in [36]
@@
As a general feature, the supercritical flow connects our calculation with the relativistic limit in both, the subluminal and in the superluminal case, because the particle production for relativistic fields happens only in the presence of a horizon. As already mentioned, the relativistic behaviour appears when the higher order of the dispersion relation can be neglected, or in other words when $\alpha\to 0$. Furthermore, when the lowest order corrections in $\alpha\ll 1$ are taken into account we saw from (17) how the relativistic Hawking temperature gets generically corrected. Now we explore such corrections for our specific dispersion relation and flow further; in doing so, we extend our investigation to a broader range of $\alpha$.
@@ IV.1 Superluminal dispersion relation
Figure 4: On the left: shape of $x_{\textsc{efh}}(\alpha)$ for the profile (28) up to $\alpha=1/2$. On the right: shape of the ratio $\kappa(\alpha)/\kappa_{\textsc{kh}}$. We see that, since $|x_{\textsc{efh}}(\alpha)|\geq|x_{\textsc{kh}}|$, so the EFH is inside the KH, the temperature of it appears to be slightly hotter than $T_{\rm H}$ at low $\alpha$ and then showing a $O(5\%)$ deviation from thermality for $\alpha\sim 0.5$
@@ IV.1.1 Low-energy regime
Here, we connect our result with the general formula provided in Eq. (17), by focusing on the low energy regime where an analytical treatment is possible. Let us start by solving (25) perturbatively in $\alpha$, one obtains
@@
The result agrees with (17), and a similar qualitative behaviour was predicted at the Killing horizon of a Hořava black hole in [27]. We find the first corrections to arise at $O(\alpha^{2})$, highlighting how the emission still remains quasi-thermal for low-energy particles. The coefficient in front of the correction is different from the one found in the black hole case in [27], however, this is not surprising because its value depends crucially on the specific geometry, in particular on the shape of $c_{s}(x)$ and $v(x)$.
@@ IV.2 Subluminal dispersion relation
For what regards the subluminal case, $\xi=-1$, we have $|c_{g}^{\rm reg}|\leq 1$ always, as such, (25) admits possible solutions only outside the Killing horizon (i.e. for positive $x$), where $|v|\leq 1$. Again, the equality is valid for $\alpha=0$, but for values close to this minimum, perturbative analyses around the Killing horizon can be safely performed. For what concerns the upper bound in the $\alpha$-range we can scrutinise (12) evaluated at $v=0$. The solutions are given by (18) with $\xi=-1$. As anticipated, we find real solutions exclusively for $\alpha\leq 1/2$. Once again, we compute the position of $x_{\textsc{efh}}(\alpha)$ as well as the value of the ratio $\kappa(\alpha)/\kappa_{\textsc{kh}}$ numerically so to test the robustness of Hawking radiation. The results are collected in Fig. 5 for the allowed range. We can see again that for $\alpha\ll 0.1$ the spectrum is basically thermal with small deviations from the relativistic result (in agreement with our previous $\alpha\to 0$ analysis in Sec. III.1.

### v2

@@ I Introduction
Here, to address these shortcomings, we employ a quasi-local technique: the tunneling picture for Hawking radiation [20, 21]. This method uses a geometric optics approximation to describe the particle creation process as a quantum tunneling event, involving complex paths across the causal boundary. In contrast to the Bogoliubov coefficients, no knowledge of the system’s asymptotic boundary is necessary, allowing particle creation to be studied almost entirely locally, wherever the event may happen in spacetime. This property is especially useful for localising effects involving effective horizons, such as those associated with modified dispersion relations that characterise fields in analogue spacetimes. Still, the tunneling method has its own limitations such as the local nature of its insights, which remain oblivion to propagation effects (e.g. gray body factors), or the intrinsic assumption of the validity of the WKB approximation (something shared with the Bogoliubov based calculations and well established at the horizon also in presence of modified dispersion relations).
@@
Note, the scope is of this paper is to be a primer into the tunneling formalism for analogue systems and to illustrate its usefulness in capturing the salient features of the Hawking effect in acoustic flows. As such, we will limit ourselves to the process of local particle production, while a full agreement with the complexity of the analysis of the Bogoliubov coefficients would require additional steps that go beyond the aim of this work.
@@ II General setting
where $c_{s}(x)$ describes the speed of sound $c^{2}_{s}\equiv\partial p/\partial\rho$, while $v(x)$ represents the velocity profile of the fluid [5]. Normally, in analogue gravity experiments, sonic horizons are realised by keeping $c_{s}$ constant and varying the velocity profile or vice versa (for example, by a Feshbach resonance in a Bose–Einstein condensate). In what follows, we shall adopt for simplicity (and easier transposition to Lorentz breaking spacetime settings) the first ansatz and assume $c_{s}={\rm constant}$.
In (1), the sign of $v(x)$ determines the nature of this spacetime. This line element has the same structure as the Schwarzschild one in Painlev’e-Gullstrand coordinates, which covers a black hole region or a white hole one depending on if $v(x)\leq 0$ or $v(x)\geq 0$, respectively [23]. This is why, as in most of the extant literature, we shall adopt the convention of a negative velocity flow transitioning from subsonic to supersonic as $x$ goes from positive to negative, i.e. with a black hole exterior on the r.h.s. of the $x$-axis). Note also that hereinafter we will omit to write the $x-$dependence explicitly, unless necessary.
We now notice that this geometry enjoys a Killing vector, associated to time translation invariance in the lab frame, given by $\chi^{a}\partial_{a}=\partial_{t}$. Together with such a lab frame, it will be useful to introduce the so-called preferred frame, which is the system of reference where the fluid is at rest, described by the two normalised vectors
@@ II.1 Perturbations with modified dispersion relation
In acoustic gravity, long wavelength perturbations, the so-called phonons, behave like a massless scalar field $\phi$ on the background (1) (see e.g. [24]). They obey, therefore, the Klein-Gordon equation $\Box\phi=0$ and enjoy a linear dispersion relation $\omega^{2}(k)=c^{2}_{s}k^{2}$. However, at short wavelengths the microstructure of the effective spacetime, characterised by some scale $\Lambda$, emerges from more general dispersion relations of the form $\omega^{2}(k)=c^{2}_{s}F(k^{2},\Lambda)$ (where we are assuming invariance under parity and that for low $k$ one has $F(k^{2},\Lambda)=k^{2}+{O}(k^{4},\Lambda)$ to recover the relativistic limit). Hence, the equation of motion for the perturbations becomes [25]
$-(\partial_{t}+\partial_{x}v)(\partial_{t}+v\partial_{x})\phi+c^{2}_{{s}}F(\partial_{x}^{2},\Lambda)\phi=0.$ (4)
This setup accommodates two free functions: the fluid’s velocity profile $v(x)$ and the dispersion relation function $F(k^{2},\Lambda)$. For what concerns our stationary flow’s velocity profile $v(x)$, we shall assume that this quantity can in general be described by a monotonous function and we will treat setups in which the flow admits a subsonic and a supersonic region (super-critical flow) or a purely subsonic setup (sub-critical flow) separately. [^1] Later on, we shall also discuss purely supersonic flows, but only as a dual case with respect to the purely subsonic ones. We address the limiting case in which the flow is everywhere sub-critical except for a sonic point (critical flow) too.
For what concerns the dispersion function, $F(k^{2},\Lambda)$ we limit ourselves to the first order corrections to the phononic dispersion relation, i.e. to modified dispersion relations of the form $\omega^{2}=c^{2}_{s}\left(k^{2}+{O}(k^{4},\Lambda)\right)$. Such dispersion relations could stem from a modified Klein-Gordon equation of the form
@@
In what follows, we shall call the dispersion relations with $\xi=\pm 1$ superluminal (upper sign) and subluminal (lower sign) respectively, as they correspond to cases for which the group velocity of perturbations is always larger/smaller than the speed of sound $c_{s}$. This is somewhat an abuse of language as these scenarios would be more precisely labelled as supersonic/subsonic but we decided to avoid such labelling to not risk any confusion with the nature of the underlying hydrodynamic flow.
@@ II.2 Particles
The trajectory of the associated ray is determined by minimising the classical action. Particles on this curve will move with the group velocity determined from the dispersion relation of the system. Introducing the four-momentum $k_{a}$ enables us to rewrite the field equation as a dispersion relation for a point particle, at the leading order in the WKB formalism [^2] For the sake of simplicity, we limit our analysis to dispersion relations without a deep minimum in the ultraviolet. Indeed, such minima are known to cause roton instabilities that are in tension with the universality of the Hawking effect and deserve a more situational analysis of the particle production, cf. [26, 27].
$\omega^{2}=c_{s}^{2}\left(k^{2}+\xi\frac{k^{4}}{\Lambda^{2}}\right)\,.$ (7)
@@
Since $\Omega$ is associated to a translational symmetry, the idea is to find mode solutions $\phi_{\Omega}$ of (7) at fixed Killing energy, which can be proven to be a conserved quantity even for a modified dispersion [28]. This, together with the dependence of our modified dispersion relation on the effective field theory scale $\Lambda$, strongly suggest the introduction of a dimensionless parameter $\alpha\equiv\Omega/\Lambda$ that captures most of the mild deviations from the relativistic behaviour (recovered in the limit $\alpha\to 0$ if one excludes the special case of the zero-mode).
@@ II.3 Particle trajectories on the background
The boundary between these two regions depends on the energy and is located at the point $x_{\textsc{tp}}(\alpha)$ (with TP standing for “turning point”) where two of the four solutions become degenerate. In terms of trajectories, this represents two smoothly merging trajectories at $x_{\textsc{tp}}(\alpha)$. The name turning point can be understood graphically in the $(t,x)$-plane. The curves at the meeting point can always be interpreted as two branches of a single trajectory that, in the $(t,x)$-plane, turns back at $x_{\textsc{tp}}(\alpha)$. Additionally, the position of $x_{\textsc{tp}}(\alpha)$ depends crucially on the sign of $\xi$. In the superluminal case, $x_{\textsc{tp}}(\alpha)$ can always be found inside the horizon, where $|\chi|^{2}>0$, while in the subluminal case $x_{\textsc{tp}}(\alpha)$ lies outside. This follows directly from the discriminant of (12). The shape of the trajectories is sketched for both cases in Fig. 1.
Figure 1: Numerical repesentation of the four different solution of (12) at fixed $\Omega$. The horizontal axis shows $v$, without specifying any profile $v(x)$ yet, while $c_{s}=1$ everywhere. The dashed black vertical line is the Killing horizon $v=-1$ in both figures. In all plots, we have taken $\alpha=0.02$ for each ray. Left panel: subluminal case ($\xi=-1$); we see that the turning point, where the dashed and the solid orange lines meet, lies outside the Killing horizon. Right panel: superluminal case ($\xi=1$) for which we find turning point to be inside the Killing horizon. Both cases share a regular mode (in blue) which travels inwards and a mode (in red) which lingers at the horizon. The latter changes its direction depending on the sign of $\xi$, while the blue one remains qualitatively unchanged.
@@ III Hawking radiation by tunnel method
In relativistic settings, i.e.  when considering phononic dispersion relations, the presence of a sonic/Killing horizon is sufficient for sonic black holes to emit particles. Essentially, such a process reflects the analogue of the standard Hawking effect [30]. Within a tunneling approach [20, 21], this correspondence can easily be seen by analysing the outgoing particle peeling off outwards from the horizon and its counterpart peeling off inwards on the other side. This approach has been substantiated by several investigations, e.g. [31] where its connection with the quantum correlation function across the horizon is shown by means of algebraic quantum field theory.
Let us stress, that while original derivation through null-geodesics [20] relied on specific coordinates and had to include the effect on the horizon of the Hawking quanta emission — so mixing the cause-and-effect relation between Hawking radiation and the shrinking of the horizon — here we shall adopt an Hamilton–Jacobi formalism providing a physically more sensible framework which does not require to include such backreaction. [^3] It is interesting nevertheless to note, that in some analog systems like in a canonical acoustic black hole in Bose-Einstein condensates, the Hawking radiation backreaction can been studied — simply it is ruled by geometrodynamical equations different from the Einstein ones — and it was found to lead to a shrinking of the horizon as in general relativity  [32]. This is not so surprising: for example it is what one would expect for an idealized radial flow in a sink endowed with an acoustic horizon. For a regular isolated system, the energy extracted via the related Hawking process can only be provided by the flow kinetic energy. The Hawking process would then induce a reduction of the flow speed (a change in the velocity profile) implying the realization of the sonic horizon deeper down in the sink geometry, i.e. a shrinking of the acoustic black hole. For a more detailed discussion, and a proof of the equivalence of the two methods, see [33].
If $x_{\textsc{kh}}$ be the horizon’s location, the tunneling rate will be given by [33, 34]
@@
It is important to note that, at least locally, the tunnelling method and the Bogoliubov coefficient approach are equivalent (cf. Appendix A). However, when considering asymptotic fluxes, the Bogoliubov method is more comprehensive as it accounts for all propagation-related effects. In this sense, our present work should be viewed as a pedagogical model for capturing the basic mechanisms of particle production, rather than a method intended to match experimental data in realistic scenarios.
III.1 Non-relativistic case
Mathematically speaking, in the above relativistic treatment, non-zero transition rates presume the presence of a simple pole in the expression of $\frac{\mathrm{d}t}{\mathrm{d}x}$. However, in a non relativistic scenario, nothing comparable will ever occur, given that the modified dispersion relation, never produces an exact peeling (see Fig. 1) and therefore the particle trajectories at the Killing horizon remain analytic. In turn, one would expect an absence of Hawking radiation in this settings (or, at least, that this effect cannot be described via the tunneling method).
@@ III.1 Non-relativistic case
In the following section, we shall focus on how to describe this effect in a simple, analytical way via the tunneling method. To this aim, we assume the validity of the Hamilton-Jacobi approach in the case of dispersive modes [^4] An extension of the Hamiltonian geometry of the phase space to particles with modified dispersion relations can be found in [39].. As we shall see, the extension of the latter to setups with modified dispersion relations will prove itself useful and informative, thus, nicely complementary to previous, semi-analytical or numerical studies  [12, 13, 14, 15, 16, 17, 18, 40, 19, 37], based on the Bogoliubov coefficients analysis.
@@ IV Particle production: supercritical flow
Our starting point will be the supercritical flow, that is to say, the permitted range for $v(x)$ supports the presence of a horizon, i.e. $|v(x)|>1$ for some $x$. For now, we keep our treatment as general as possible without particularising to a specific profile for $v(x)$. We demand: i) $v(x)$ to be monotonous to avoid inner Killing horizons (namely, $v(x)=-1$ has a single root), ii) $v\to 0$ for $x\to-\infty$, and iii) conditions (18) and (21) hold almost everywhere apart from a small region around the effective horizon. A typical example of such flow is the one considered in [41] and shown in figure 3
@@
As a general feature, the supercritical flow connects our calculation with the relativistic limit in both, the subluminal and in the superluminal case, because the particle production for relativistic fields happens only in the presence of a horizon. As already mentioned, the relativistic behaviour appears when the higher order of the dispersion relation can be neglected, or in other words when $\alpha\to 0$. Now we explore such corrections for our specific dispersion relation and flow further; in doing so, we extend our investigation to a broader range of $\alpha$.
@@ IV.1 Superluminal dispersion relation
Figure 4: On the left: shape of $x_{\textsc{efh}}(\alpha)$ for the profile (26) up to $\alpha=1/2$. On the right: shape of the ratio $\kappa(\alpha)/\kappa_{\textsc{kh}}$. We see that, since $|x_{\textsc{efh}}(\alpha)|\geq|x_{\textsc{kh}}|$, so the EFH is inside the KH, the temperature of it appears to be slightly hotter than $T_{\rm H}$ at low $\alpha$ and then showing a $O(5\%)$ deviation from thermality for $\alpha\sim 0.5$. Note that we can directly compare the positions due to the monotony of $v(x)$
@@ IV.1.1 Low-energy regime
Here, we focus on the low energy regime where an analytical treatment is possible. Let us start by solving (23) perturbatively in $\alpha$, one obtains
@@
The result shows a similar qualitative behaviour was predicted at the Killing horizon of a Hořava black hole in [29]. We find the first corrections to arise at $O(\alpha^{2})$, highlighting how the emission still remains quasi-thermal for low-energy particles. The coefficient in front of the correction is different from the one found in the black hole case in [29], however, this is not surprising because its value depends crucially on the specific geometry, in particular on the shape of $c_{s}(x)$ and $v(x)$.
@@ IV.2 Subluminal dispersion relation
For what regards the subluminal case, $\xi=-1$, we have $|c_{g}^{\rm reg}|\leq 1$ always, as such, (23) admits possible solutions only outside the Killing horizon (i.e. for positive $x$), where $|v|\leq 1$. Again, the equality is valid for $\alpha=0$, but for values close to this minimum, perturbative analyses around the Killing horizon can be safely performed. For what concerns the upper bound in the $\alpha$-range we can scrutinise (12) evaluated at $v=0$. The solutions are given by (16) with $\xi=-1$. As anticipated, we find real solutions exclusively for $\alpha\leq 1/2$. Once again, we compute the position of $x_{\textsc{efh}}(\alpha)$ as well as the value of the ratio $\kappa(\alpha)/\kappa_{\textsc{kh}}$ numerically so to test the robustness of Hawking radiation. The results are collected in Fig. 5 for the allowed range. We can see again that for $\alpha\ll 0.1$ the spectrum is basically thermal with small deviations from the relativistic result.